a ball is thrown upward with an initial velocity of 64 ft/sec. from a height of 720 ft. its height h, in feet, after t seconds is given by h (t)=-16t^2+64t+720. After how long will the ball reach the ground?
Answer by lwsshak3(11628)

a ball is thrown upward with an initial velocity of 64 ft/sec. from a height of 720 ft. its height h, in feet, after t seconds is given by h (t)=-16t^2+64t+720. After how long will the ball reach the ground?
***
When the ball reaches the ground, its height=0
-16t^2+64t+720=0
-t^2+4t+45=0
t^2-4t-45=0
(t-9)(t+5)=0
t=9
After how long will the ball reach the ground? 9 sec